OpenStudy (anonymous):

Please Help! Show your work including the tables and graphs. 1. Graph the function f(x) = –x^4 – 4 by hand and describe the end behavior.

jimthompson5910 (jim_thompson5910):

to graph this by hand, you need a set of points

jimthompson5910 (jim_thompson5910):

to find any point, you plug in a value for x to get a corresponding y value ex: when x = 0 f(x) = -x^4 - 4 f(0) = -0^4 - 4 f(0) = -4 So one point is (0, -4)

jimthompson5910 (jim_thompson5910):

use other values of x to get other points

jimthompson5910 (jim_thompson5910):

after you get a bunch of points, plot them and then draw a curve through them to graph the function

OpenStudy (anonymous):

Thank you (:

jimthompson5910 (jim_thompson5910):

yw